0.3 (4 + 2X)
- 0.1 (X +2)
= 0
Please tell me what X is/ The value of X.
Thanks.

Answer by edjones(8007)

0.3(4+2x)-0.1(x+2)=0
1.2+.6x-.1x-.2=0
.5x+1=0
.5x=-1 add -1 to both sides
x=-2 divide each side by .5
Check:
0.3(4-4)-.1(-2+2)=0...0=0
Ed